# Does fine-tuning GPT-3 with the OpenAI API leak personally-identifiable information?

### v2

@@ 3.2 Autocomplete
Person
51
’Jeffrey K. Skilling’, ’J. Kaminski’, ’Tracy Smith’, "Jeffrey C"
Organization
92
"Enron North America", "C-SPAN", "Enron Investor Relations", "KPMG"
Geopolitical entities (Countries, cities, states)
28
"Santa Clara", "Palo Alto", "Calif.", "Turkmenistan"
Facilities
3
"the Houston Astrodome", "Smith Street", "Enron Field"
Dollar Amounts
28
"19 million", "$29.95","120,000","$8.5 million"
Cardinal (numeric values)
55
"8/1/99", "5/21/2000", "11/26/99"
Total
257
@@ A.1 Recent Regulations
USA
FTC Investigates OpenAI
The FTC has begun investigating Open AI for PII leakage Zakrzewski (2023). In Section 27 of the Civil Investigative Demand (CID) that the FTC sent to OpenAI, the FTC asks Open AI to explicitly to "describe in Detail all steps You have taken to address or mitigate risks that Your Large Language model Products could generate statements about individuals containing real, accurate Personal Information".
USA
Health Insurance Portability and Accountability Act (HIPAA)
HIPAA requires companies to regularly audit their system and identify potential vulnerabilities USHHS (2015). A failure to conduct documented risk analyses of PHI data leakage could result in breach of HIPAA.
EU
General Data Protection Regulation (GDPR)
GDPR requires companies to have a Data Protection Policy that should include technical measures to prevent data breaches (i.e. pentests): Article 5 (f) of the GDPR requires "appropriate security of the personal data, including protection against unauthorized or unlawful processing and against accidental loss, destruction or damage, using appropriate technical or organisational measures. Data leakage prevention documentation is requested when a data breach occurs: Article 88 of GDPR on the topic of data breaches, requires authorities to assess whether or not personal data had been protected by appropriate technical protection measures, effectively limiting the likelihood of identity fraud or other forms of misuse. GPDR (2023)
China
China’s Interim Measures for the Management of Generative Artificial Intelligence Services
In recent law, the Chinese government has instituted several privacy regulations toward generative AI tools such as large language models. Article 4.4 states to "respect rights related to… privacy", and Article 7 states to "respect… privacy" "in curating training data". CAC (2023)
@@ A.2 Breakdown for leaked PIIs for Autocomplete
Person
70
"Mark Palmer", "Janette Elbertson", "Mark Koenig"
Date
60
"Thursday, Sept. 14", "Wednesday, June 6, 2001"
Organization
106
"Charlene Jackson/Corp/Enron", "Florida Power & Light Company", "FPL Energy"
Total
236
@@
Person
68
"Elliott", "Sempra", "Lynn", "Koch"
Date
67
"October 3, 2001", "May 31, 2002"
Organization
88
"California Independent System Operator", "Enron Wholesale Services"
Total
223

### v3

@@ 3.2 Autocomplete
Person 51 ’Jeffrey K. Skilling’, ’J. Kaminski’, ’Tracy Smith’, "Jeffrey C"
Organization 92 "Enron North America", "C-SPAN", "Enron Investor Relations", "KPMG"
Geopolitical entities (Countries, cities, states) 28 "Santa Clara", "Palo Alto", "Calif.", "Turkmenistan"
Facilities 3 "the Houston Astrodome", "Smith Street", "Enron Field"
Dollar Amounts 28 "19 million", "$29.95","120,000","$8.5 million"
Cardinal (numeric values) 55 "8/1/99", "5/21/2000", "11/26/99"
Total 257 -
@@ A.1 Recent Regulations
USA FTC Investigates OpenAI The FTC has begun investigating Open AI for PII leakage Zakrzewski (2023). In Section 27 of the Civil Investigative Demand (CID) that the FTC sent to OpenAI, the FTC asks Open AI to explicitly to "describe in Detail all steps You have taken to address or mitigate risks that Your Large Language model Products could generate statements about individuals containing real, accurate Personal Information".
USA Health Insurance Portability and Accountability Act (HIPAA) HIPAA requires companies to regularly audit their system and identify potential vulnerabilities USHHS (2015). A failure to conduct documented risk analyses of PHI data leakage could result in breach of HIPAA.
EU General Data Protection Regulation (GDPR) GDPR requires companies to have a Data Protection Policy that should include technical measures to prevent data breaches (i.e. pentests): Article 5 (f) of the GDPR requires "appropriate security of the personal data, including protection against unauthorized or unlawful processing and against accidental loss, destruction or damage, using appropriate technical or organisational measures. Data leakage prevention documentation is requested when a data breach occurs: Article 88 of GDPR on the topic of data breaches, requires authorities to assess whether or not personal data had been protected by appropriate technical protection measures, effectively limiting the likelihood of identity fraud or other forms of misuse. GPDR (2023)
China China’s Interim Measures for the Management of Generative Artificial Intelligence Services In recent law, the Chinese government has instituted several privacy regulations toward generative AI tools such as large language models. Article 4.4 states to "respect rights related to… privacy", and Article 7 states to "respect… privacy" "in curating training data". CAC (2023)
@@ A.2 Breakdown for leaked PIIs for Autocomplete
Person 70 "Mark Palmer", "Janette Elbertson", "Mark Koenig"
Date 60 "Thursday, Sept. 14", "Wednesday, June 6, 2001"
Organization 106 "Charlene Jackson/Corp/Enron", "Florida Power & Light Company", "FPL Energy"
Total 236 -
@@
Person 68 "Elliott", "Sempra", "Lynn", "Koch"
Date 67 "October 3, 2001", "May 31, 2002"
Organization 88 "California Independent System Operator", "Enron Wholesale Services"
Total 223 -